A 12µF capacitor is required to store 4J of energy. Find the p.d to which the capacitor must be charged.
1
Expert's answer
2020-04-28T09:45:34-0400

"E=4" J

"C=12\times 10^{-12} \mu F"

"E={\frac {CU^2} 2}"

"U=\sqrt{{\frac {2E} C}}=8.2\times10^5 V"